OpenStudy (anonymous):

On an architect's blueprint, 1 inch corresponds to 6 feet. If an exterior wall is 26 feet long, find how long the blueprint measurement should be. Write the answer as a mixed number if necessary.

OpenStudy (anonymous):

1inch=0.083ft 1inch:6ft \[26/6=13/3=4(1/3)\]

OpenStudy (anonymous):

i wasn't sure how to write the mixed number. just note the 4 1/3 is the mixed number.

OpenStudy (anonymous):

and not to be multiplied.